40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Sat Oct 26 21:46:33 2013, RANA, Update, IOO, Mode Cleaner Tune-UP 
|
Mon Oct 28 14:12:48 2013, manasa, Update, IOO, Mode Cleaner relocked
|
Wed Oct 30 01:53:52 2013, Jenne, Update, IOO, Mode Cleaner relocked
|
Mon Oct 28 21:33:55 2013, RANA, Update, IOO, Mode Cleaner Tune-UP
|
Thu Oct 31 15:34:28 2013, manasa, Update, IOO, MC not happy since yesterday
|
Thu Oct 31 20:05:48 2013, RANA, Update, IOO, Mode Cleaner Tune-UP
|
Fri Nov 1 09:45:32 2013, Steve, Update, IOO, beam dumps to be find
|
Mon Nov 4 11:37:12 2013, Steve, Update, IOO, PSL output shutter installed
|
Mon Nov 4 12:59:43 2013, Jenne, Update, IOO, MC alignment not so good after PSL output shutter installed
|
Tue Nov 5 00:39:43 2013, manasa, Update, IOO, IFO alignment tuning
|
Tue Nov 5 08:44:21 2013, Steve, Update, IOO, after last steering mirror mount swap 
|
Tue Nov 5 17:12:34 2013, Steve, Update, LSC, centered qpds
|
Wed Nov 6 08:33:53 2013, Steve, Update, IOO, Poiting changes of PSL 
|
Thu Nov 7 08:57:20 2013, Steve, Update, IOO, PSL pointing monitoring 
|
Mon Nov 11 22:35:45 2013, RANA, Update, IOO, PSL pointing monitoring
|
Tue Nov 12 23:48:23 2013, RANA, Update, IOO, PSL pointing monitoring
|
Thu Nov 21 16:32:20 2013, Steve, Update, IOO, PMC needs a touch of love
|
Fri Dec 6 15:57:41 2013, Steve, Update, IOO, beam dumps for PSL pointing monitoring 
|
Tue Dec 10 10:07:01 2013, Steve, Update, IOO, more beam traps 
|
|
Message ID: 9421
Entry time: Thu Nov 21 16:32:20 2013
In reply to: 9365
|
Author: |
Steve |
Type: |
Update |
Category: |
IOO |
Subject: |
PMC needs a touch of love |
|
|
The PMC power degrading on this 3 days plot. MC2 -T = 14,200 counts. C1:IOO-MC_TRANS_SUM can not be ploted in dataviewer. The MEDM screen has a valid number.
Initial pointing is not so bad (what does "not so bad" mean ???)
C1iscey comes and goes again.
|
|
|